OpenStudy (anonymous):

Translate the following verbal statement into an algebraic equation and then solve: Use x for your variable. The difference of two times a number and five is eight.

OpenStudy (anonymous):

2*x-5=8

OpenStudy (anonymous):

so it would be 2x-5=8 +5 2x=13 13/2 x=6.5

OpenStudy (defiance):

6.5 is right, so yes.

OpenStudy (defiance):

Cause if you reverse it, you can do 8+5 = 13 13 divided by 2 = 6.5

OpenStudy (defiance):

Your equation is right aswell.